Magic Realm, Inc., has developed a new fantasy board game. The company sold 48,000 games last year at a selling price of $67 per game. Fixed expenses associated with the game total $864,000 per year, and variable expenses are $47 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor.

Required:

1-a. Prepare a contribution format income statement for the game last year.

1-b. Compute the degree of operating leverage.

2. Management is confident that the company can sell 61,440 games next year (an increase of 13,440 games, or 28%, over last year). Given this assumption:

a. What is the expected percentage increase in net operating income for next year?

b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

1a) Contribution income statement

Total Per unit
Sales 3216000 67
Variable cost 2256000 47
Contribution margin 960000 20
Fixed cost 864000
Operating income 96000

1b) Degree of operating leverage = Contribution margin / operating income = 960000/96000 = 10

2a) Net operating income increase by 28*10 = 280%

2b) Expected operating income = 96000*3.8 = $364800